OpenStudy (anonymous):

A trapezoid has an area of 75 square inches and it's two bases are 8 and 17 inches long What's the height of the trapezoid?

OpenStudy (campbell_st):

well you know \[75= \frac{h}{2} \times( 8 + 17)\] now make h the subject... to find the height

OpenStudy (anonymous):

Still confused

OpenStudy (campbell_st):

ok... what about this after simplifying you get \[75 = \frac{h}{2} \times (25)\] any thoughts...?

OpenStudy (anonymous):

No😳

OpenStudy (campbell_st):

ok... multiply both sides of the equation by 2 then divide both sides of the equation by 25 this will allow you to find the height

OpenStudy (anonymous):

I'm just confused as to which numbers I'm multiplying by 25?

OpenStudy (campbell_st):

mutiply 75 by 2 then divide by 25 the answer will be the height of the trapezoid

OpenStudy (anonymous):

So is the answer 6

OpenStudy (campbell_st):

thats it....